OpenStudy (anonymous):

Factor completely 1. a^3-ab+a^2-b 2. g^5-g

OpenStudy (anonymous):

\[ a^3-ab+a^2-b = (1+a) \left(a^2-b\right) \] \[g^5-g= (g-1) g (1+g) \left(1+g^2\right) \]

OpenStudy (anonymous):

Can you do it in steps or explains?